OpenStudy (anonymous):

@InYourHead @inkyvoyd @irishchelsc @iamanocean @some_someone @stabar @SnuggieLad @Will679 @dusty can someone help me i dont want the answers i want to learn

OpenStudy (anonymous):

the first question is x^2+20x+36

OpenStudy (anonymous):

um i suck at math basically so idk so sry bb </3

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

note (x+a)(x+b) - this will be the form of your solution (x+a)(x+b)=x^2+(a+b)x+ab = x^2+20x+36

OpenStudy (anonymous):

so given that, we can conclude that a+b=20 a*b= 36 then solve for a and b

OpenStudy (anonymous):

u just lost me

OpenStudy (anonymous):

Can I step in for a moment? Let's look at our problem here.... \[x^{2}+20x+36\] You see that number? It's 36. And see that middle number? It's 20. So, we have to choose two numbers .... that MULTIPLY to make 36, and ADD TOGETHER to make 20. Can you think of two numbers that would fit both of those rules?

OpenStudy (anonymous):

(x+a)(x+b) when you foil that, multiply it out, etc you will get x^2+ (a+b) x +a*b

OpenStudy (anonymous):

That LAST number is 36. I'm sorry, I forgot to mention "last."

OpenStudy (anonymous):

\[x^2+20x+36\] \[(x+2) (x+18)\]

OpenStudy (anonymous):

6X6????

OpenStudy (anonymous):

Good. some_someone figured it out. The numbers are 18 and 2. 18 times 2 is 36. 18 plus 2 is 20. You get it?

OpenStudy (anonymous):

6 times 6 IS 6. But 6 + 6 = 12. Not 20. Good try though.

OpenStudy (anonymous):

sorry i didnt know u had to use it for both ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Once we figure out the two numbers, we write it out like this: (x + 18) (x + 2)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

This was a very easy one problem. Do you have anymore like it? If not, I can give you another example, and we can see if you can solve it on your own.

OpenStudy (anonymous):

Sorry about the typo.

OpenStudy (anonymous):

try 20 more lol

OpenStudy (anonymous):

Damn.

OpenStudy (anonymous):

Okay. Give us another.

OpenStudy (anonymous):

the next one is factor the polynomial 8x^4+6x^3-2x^2

OpenStudy (anonymous):

(reverse) distributive property of multiplication what common multiple do they all share?

OpenStudy (anonymous):

note they're all even they all have variable x

OpenStudy (anonymous):

24

OpenStudy (anonymous):

sorry i meant common factor

OpenStudy (anonymous):

@completeidiot Sorry, I'm having trouble with this one. Is the common factor \[2x^{2}\] ?

OpenStudy (anonymous):

yes, you can factor out a \[2x^2\] \[8x^4+6x^3-2x^2=2x^2(4x^2+3x-1)\]

OpenStudy (anonymous):

what's left is factoring 4x^2+3x-1

OpenStudy (anonymous):

I solved it. Thank you very much.

OpenStudy (anonymous):

@danya1 Let's leave the \[2x^{2}\] to the side. Forget about that part, for now. We've broken it down to this.... \[4^{2}+3x-1\]

OpenStudy (anonymous):

We have to factor that part now. Do you know PEMDAS?

OpenStudy (anonymous):

Sorry. FOIL.

OpenStudy (anonymous):

You know... First Outside, Inside Last You know about that?